Grade 11IIT JEE Entrance Exam

A copper wire (Y=1*10^11N/m2) of length 6m and a steel wire (2*10^11N/m2) of length 4m each of cross section 10^-5m2 are fastened end to end A load of 100N is applied to one end keeping the other end fixed the elongation produced in the copper wire is:
a)0.2mm b)0.4mm c)0.6mm d)0.8mm

Pratima  chakraborty
6 Years agoGrade 11

2 Answers

Bhargav Petkar
6 Years ago
YCU=1011N/m2                                          YSTEEL=2*1011N/m2
lCu=6m                                                        lsteel=4m
therefore,change in Y=(2*1011 - 1011)  N/m2    = 1011N/m2
change in length     = 2m
hence , elongation =FL/AY=100*2/10-5*1011=0.2mm